# LLM4SCREENLIT: Recommendations on Assessing the Performance of Large Language Models for Screening Literature in Systematic Reviews

### v1

@@ 3.4 Reporting Variation and Confidence Intervals
The set of papers we reviewed included six papers that reported measures of variance or confidence intervals, see Figure˜3. However, when we have true and predicted classification for all abstracts related to a specific SR, we have the complete population not a sample, so the concept of sampling error is meaningless. Thus, if we report confidence intervals, we need to be very clear to which population and experimental hypotheses they apply.
One situation where CIs are important is when researchers intend to use results of a validation exercise based on a random subset of the available abstracts to decide whether an LLM can be used to analyse the remaining abstracts. In this case, we need to calculate the mean and variance of appropriate performance metrics from the validation sample results. For this purpose, we would suggest using multiple resampling of the validation sample without replacement, see the example visualisation in Figure˜4 produced by our R simulation script, in which we reused the known performance metrics of the two models (Model A: llama3-Athene:70b, and Model B: llama3.1.:8b) reported in Table˜1222We developed an R script that generated 10000 subsample distributions of different sizes (100, 200, 300, 400, 500) for MCC and WMCC that are in line with the performance metrics reported in Table 1 and visualized the results.. Such visualisations may help to decide whether LLMs can be used to analyse the remaining abstracts, and which models to choose.
@@
It is also worth mentioning, that we do not recommend CIs based on the binomial distribution because this assumes that all the data items are independently and identically distributed (iid), which is unlikely for a set of abstracts, and, for the same reason, we do we recommend bootstrapping (i.e., sampling with replacement).
@@ 3.5 Review Limitations
A limitation of this review is that it is based on a convenience sample of primary studies. We sourced papers from our own knowledge, but in order to address the criticism that we might have explicitly selected papers that support our own opinions, we also included all the primary studies included in two systematic reviews. Furthermore, the problems we found in the DC+ paper were also found in papers from all three sources. Likewise, good practices varied across papers from different sources. It is perhaps surprising how little overlap there was between the three data sources. This would suggest that there are potentially many more papers on the topic than the 28 used in this review.
@@ 4 Discussion and Conclusions
In this paper, we have demonstrated the critical challenges associated with evaluating LLMs for literature screening in SRs. We have used the recent evaluation study DC+ [4] as a means of illustrating some of the challenges of choosing appropriate classification metrics. DC+ authors report and discuss thoroughly misleading metrics such as Accuracy, which suggest poorly performing tools are to be preferred. They also omit any sense of cost or the idea that FPs are less problematic than FNs. Notably, across three SR datasets and 18 LLMs in DC+, screening performance was frequently undermined by substantial Lost Evidence (i.e., high FN rates), see Figure˜1, demonstrating that correctness/accuracy and related metrics can be seriously misleading under class imbalance and asymmetric misclassification costs. Our review of practices in 27 other papers confirms that these problems are not unique to DC+.
We believe performance metrics problems arise because researchers do not always select metrics that are mapped to the actual needs of the problem domain. A better approach is to report the individual confusion matrix counts together with a small number of performance metrics focussed on the specific research questions being investigated. This means that evaluations of literature screening must prioritize chance-anchored metrics such as MCC and, additionally, explicitly reflect FN vs. FP asymmetry via cost-sensitive analysis. For this purpose, we propose Weighted MCC (WMCC) as a principled extension that retains MCC’s correlation meaning and robustness to class imbalance, while addressing the challenge of asymmetric misclassification costs by encoding domain-specific cost ratios.
@@ 4.1 Implications
There are many implications stemming from this paper for researchers and practitioners, as well as for those responsible for setting journal and conference policies. Researchers should prioritize prospective, leakage-aware benchmarks and standardize WMCC reporting to enable credible conclusions and robust meta-analytic synthesis across SRs. Practitioners should adopt a reporting kit of their LLM validation exercises that includes complete confusion matrices, Lost Evidence, and WMCC with confidence intervals. They could also set in advance a maximum acceptable Lost Evidence (minimum acceptable recall) threshold that reflects the risk tolerance and objectives of the specific review (they would likely be less stringent for software engineering SRs than for high‑stakes clinical SRs), so that Lost Evidence from FNs is bounded by design. In practice, this means that if preliminary validation of LLMs suggests that LLM performance cannot meet or exceed acceptable levels for a specific SR task, and there are no clear indications of how prompts could be usefully refined, the task must be performed by human researchers
Journals, conferences, and SR guidelines should require confusion matrices, uncertainty estimates, baseline comparators, explicit leakage/contamination statements, and open artifacts (prompts, seeds, and any materials/artifacts), while discouraging accuracy-focused reporting.
@@ 4.2 Recommendations
Although, in Sections˜2.5 and 3.2, we have discussed a variety of good practices for evaluating LLM screening performance, our main goal was to provide actionable recommendations for (i) Researchers and practitioners, as well as (ii) Policymakers (e.g., journals, conferences, guideline authors) on how to deal with the observed challenges associated with evaluating the performance of LLMs for screening literature for SRs. In addition to focusing our recommendations on specific target audiences, we decided to organize recommendations into decision-centric themes to improve comprehension. As a result, we have the following recommendations organized by target audience and themes:
@@
Target Audience: Researchers and Practitioners Theme: Metrics and cost-sensitive evaluation (R1) Standardize reporting on Lost Evidence (Recall), MCC, and Weighted MCC (WMCC) with explicit justification of FN:FP cost ratios, and report only relevant metrics while avoiding Accuracy/PABAK as primary metrics (origin: (P9) in Section˜2.5 & Section˜2.3). (R2) Base comparative conclusions on cost-sensitive analyses that reflect asymmetric misclassification costs, using WMCC to combine chance-correction with cost asymmetry and avoiding over-optimizing Recall alone (origin: (P9) in Section˜2.5). (R3111An asterisk ’*’ means that the recommendation is optional.) Predefine acceptable Lost Evidence (minimum Recall) thresholds as guardrails for the review’s risk tolerance and objectives, aligned to review type (e.g., SR, Mapping/Scoping Study, Rapid Reviews) and domain (e.g., healthcare, software engineering) (origin: Section˜4.1). Theme: Reporting and transparency (R4) Publish complete confusion matrices for every model, dataset, and prompt to enable recomputation of necessary metrics like Lost Evidence, MCC, WMCC, and alternative (e.g., cost-benefit) analyses and future meta-analyses [^3] DC+ revealed a consistent problem with lost evidence across three reviews and 18 LLMs. This important result is only clear because they reported all their confusion matrices. (origin: (P1) in Section˜2.5). (R5) For validation tests, report uncertainty for each performance metric via confidence intervals and document the estimation method used. When testing LLMs on a random sample of abstracts to decide whether to deploy them on the remaining abstracts, use resampling without replacement to estimate confidence intervals for likely performance on the full dataset (origin: (P3) in Section˜3.2; resampling method proposed by the authors in Section˜3.4 and illustrated in Figure˜4). (R6) Quantify LLM output consistency and null or invalid outputs, specify the evaluations rule that treats unclassifiable or referred-back items for fair metric computation (origin: (P4) and (P5) in Section˜3.2). (R7) Release open artifacts, including prompts, seeds, code, and curated data, as well as include non-LLM baselines for fair comparison and to support open science (origin: (P2) in Section˜2.5). Theme: Study design and validity (R8) Use prospective or temporally safeguarded evaluations and explicitly state contamination/leakage risks and mitigations to prevent training-test overlap (origin: (P7) & (P8) in Section˜3.2). (R9) Include non-LLM baselines and assess LLMs (as black-box tools) against conventional methods to ensure practical value and fair benchmarking (origin: (P2) in Section˜2.5). Theme: Decision thresholds and operations (R9) When observed Lost Evidence exceeds the pre-specified threshold, escalate to human review or adjust prompts/models to maintain SR validity (origin: Section˜4.1).

### v2

@@ 3.5 Reporting Variation and Confidence Intervals
The set of papers we reviewed included seven papers that reported measures of variance or confidence intervals, see Figure˜3. However, when we have true and predicted classification for all abstracts related to a specific SR, we have the complete population not a sample, so the concept of sampling error is meaningless. Thus, if we report confidence intervals, we need to be very clear to which population and experimental hypotheses they apply.
One situation where CIs are important is when researchers intend to use results of a validation exercise based on a random subset of the available abstracts to decide whether an LLM can be used to analyse the remaining abstracts. In this case, we need to calculate the mean and variance of appropriate performance metrics from the validation sample results. For this purpose, we would suggest using multiple resampling of the validation sample without replacement, see the example visualisation in Figure˜4 produced by our R simulation script, in which we reused the known performance metrics of the two models (Model A: llama3-Athene:70b, and Model B: llama3.1.:8b) reported in Table˜1666We developed an R script that generated 10,000 subsample distributions of different sizes (100, 200, 300, 400, 500) for MCC and WMCC that are in line with the performance metrics reported in Table 1 and visualized the results.. Such visualisations may help to decide whether LLMs can be used to analyse the remaining abstracts, and which models to choose.
@@
It is also worth mentioning that we do not recommend CIs based on the binomial distribution because this assumes that all the data items are independently and identically distributed (iid), which is unlikely for a set of abstracts, and, for the same reason, we do not recommend bootstrapping (i.e., sampling with replacement).
@@ 3.6 Review Limitations
A limitation of this investigation is that our review of existing studies is based on a convenience sample of primary studies. We sourced papers from our own knowledge, but in order to address the criticism that we might have explicitly selected papers that support our own opinions, we also included all the primary studies included in two systematic reviews [15, 16]. These SRs were the only ones we were able to find, and they were not published by SE-related journal/conferences. However, Kim et al. [15] did find one of the five SE related papers we found, so clearly SE-related papers had not been excluded. Another limitation that applies to any SR is that it includes only previously published papers and searches would have been conducted well before the SR’s publication date. However, both SRs provided selections of screening papers obtained independently of the papers we found by our informal searches. Furthermore, good and bad practices were found in papers from all three sources again suggesting that our set of papers were not biased. Although not biased, the simple fact that GenAI is a hot topic means that as soon as a set of published papers is collated and analysed, it will be out of date. This means our list of good and bad practices may be incomplete. In addition, we cannot predict the rate of adoption of various good and bad practices from the analyses presented in this paper.
@@
Finally, our recommendations were developed for title-and-abstract screening. Section˜4.3 discusses their principled extension to full-text screening and data extraction based on the non-SE evidence in Cao et al. [28] and Khraisha et al. [9], but this extension remains largely theoretical: no SE-specific empirical worked example yet exists for full-text screening or data extraction, so the transferability claims in Table˜6 should be read as hypotheses to be tested rather than validated results. Empirical validation of WMCC on SE full-text screening and on the multi-class or ordinal extraction tasks typical of SE (mapping-study classification, quantitative extraction, quality assessment) is important future work.
@@ 4 Discussion and Conclusions
In this paper, we have presented LLM4SCREENLIT—a set of recommendations for evaluating LLMs for literature screening in SRs—motivated by some critical challenges we have demonstrated. We believe performance metrics problems arise because researchers do not always select metrics that are mapped to the actual needs of the problem domain. A better approach is to report the individual confusion matrix counts together with a small number of performance metrics focussed on the specific research questions being investigated.
In particular, we propose that evaluations of literature screening must prioritize chance-anchored metrics such as MCC and, additionally, explicitly reflect FN vs. FP asymmetry via cost-sensitive analysis. For this purpose, we propose Weighted MCC (WMCC) as a principled extension that retains MCC’s correlation meaning and robustness to class imbalance, while addressing the challenge of asymmetric misclassification costs by encoding domain-specific cost ratios. We note that WMCC is not a replacement for MCC. For meta-analysis purposes, MCC allows researchers to assess the predictive capability of different LLMs across different SRs, irrespective of their choice of WMCC weight.
While MCC and WMCC summarize performance at a single operating point, we acknowledge that reporting performance across multiple operating points would strengthen evaluation studies when continuous LLM outputs (e.g., confidence scores) are available. In such cases, researchers could report Lost Evidence-workload curves, decision curve analysis, or performance at multiple Lost Evidence levels to help practitioners understand the Lost Evidence-workload trade-off. The specific levels should be determined by the domain and SR type. However, most current LLM screening studies use binary prompts (include/exclude) that naturally produce single operating points rather than tunable thresholds.
@@ 4.1 Implications
There are many implications stemming from this paper for researchers (i.e., individuals studying the capabilities of LLMs) and practitioners (i.e., individuals validating GenAI tools as part of the conduct of a specific SR), as well as for those responsible for setting journal and conference policies. Researchers should prioritize prospective, leakage-aware benchmarks and standardize WMCC reporting to enable credible conclusions and MCC for robust meta-analytic synthesis across SRs. They also need to be aware of the difference between the studies based on previously published SRs and studies undertaken to validate LLMs in the context of conducting a new SR. It is important that they address the implications of their results for practitioners in terms of how their results can support SR-based validation exercises. The implications differ by study type (see the benchmarking vs. deployment distinction in Section˜4.2). For researchers conducting retrospective benchmarks, the core requirements are complete confusion matrices, chance-anchored metrics (MCC, WMCC), and leakage safeguards. Practitioners deploying LLMs on a specific SR should adopt a reporting kit of their LLM validation exercises that includes complete confusion matrices based on a validation random sample, and performance metrics including Lost Evidence (or Recall), and MCC with confidence intervals. Thresholds for maximum Lost Evidence should be set in advance and are context, domain, and SR type dependent, reflecting the risk tolerance and objectives of the specific review, and should be determined by stakeholders based on the specific consequences of missed evidence. This means that Lost Evidence from FNs is bounded by design. Decisions can then be made based on which (if any) classification processes (i.e., specific combination of one or more LLM(s) with either one or zero human researchers) achieved acceptable values of Lost Evidence (e.g. Lower Confidence interval >0.8 and produced a genuine prediction (i.e., delivered a 95% Lower confidence value of MCC>0). Any classification processes that achieved these criteria would then be ranked based on WMCC values, with the process with the best WMCC being selected (assuming the WMCC values were greater than zero). If preliminary validation of LLMs supported classification process suggests that performance cannot meet or exceed acceptable levels for a specific SR task, and there are no clear indications of how prompts could be usefully refined, the task must be performed by human researchers.
A major outstanding issue is how the relative costs of FPs and FNs can be determined. Whether we use WMCC (or any other form cost/benefit assessment), we need a weighting factor based on the relative costs of FNs and FPs. In this study, we used 10:1 which might be acceptable for the SE domain, but we have no independent rationale for this value. For commissioned SRs, it may be possible to ask the commissioning group or other stakeholder groups for input about the issue. Assessing relative costs is also made more complex by the fact that not all papers are of equal importance, missing a study with serious methodological flaws is not as important as missing a rigorous study. This discussion also suggests that we need information about the rate of studies that are currently missed by the current gold standard process of screening by two (competent) researchers with disagreements being resolved. Such information would at least provide some information about current levels of FNs.
Journals, conferences, and SR guidelines should require confusion matrices, uncertainty estimates (when appropriate), baseline comparators, explicit leakage/contamination statements, and open artifacts (prompts, seeds, and any materials/artifacts), while discouraging accuracy-focused reporting.
@@ 4.2 Recommendations
Although, in Sections˜2.5 and 3.2, we have discussed a variety of good practices for evaluating LLM screening performance, our main goal was to provide the LLM4SCREENLIT recommendations—actionable guidance for (i) Researchers and practitioners, as well as (ii) Policymakers (e.g., journals, conferences, guideline authors) on how to deal with the observed challenges associated with evaluating the performance of LLMs for screening literature for SRs. In addition to focusing our recommendations on specific target audiences, we decided to organize recommendations into decision-centric themes to improve comprehension. To provide practical guidance, Figures˜5 and 6 present two decision trees—one for benchmarking studies and one for deployment studies—that illustrate the workflows for evaluating LLMs for SR screening. Each tree is organized into two parallel tracks: (1) Study Design steps to complete before running the evaluation, and (2) Metric Selection steps to apply during and after evaluation. These tracks converge at a cost assessment phase and a reporting phase (which explicitly includes reporting all confusion matrix elements). The benchmarking tree (Figure˜5) terminates with a reporting endpoint, while the deployment tree (Figure˜6) adds a decision point based on the predefined Lost Evidence threshold, leading to either deployment or escalation to human review.
Before presenting the recommendations, we distinguish between two study types that share most of the evaluation mechanisms but differ in their purposes: benchmarking studies and deployment studies. LLM benchmarking studies usually evaluate one or more LLMs on previously completed SRs, in order to compare models, prompts, or datasets. Their key question is “How well do LLMs perform at SR screening?LLM deployment studies are usually prospective. Typically, they test LLMs on a random sample of candidate primary studies from an SR that is currently being conducted, in order to decide whether to deploy the LLM on the remaining (unscreened) abstracts. Their key question is “Should this LLM be used to screen my SR?”. Because they lead to an operational decision, they need a pre-specified Lost Evidence threshold, confidence intervals from the sample, and an escalation rule if the threshold is not met. We tag each recommendation R1–R10 below with the study type(s) to which it applies.
BENCHMARKING — LLM EVALUATION How well do LLMs perform on pre-labelled data? No deployment decision. STUDY DESIGNMETRIC SELECTION Temporal safeguards / leakage statement (R8) Non-LLM baselines (R9) only if the benchmark aims to inform SR practice and labels are available (e.g., SR update) Compute confusion matrix (R4) Calculate Lost Evidence/Recall (R1) Calculate MCC (R1) COST ASYMMETRY ASSESSMENT Explore range of FN:FP cost ratios w (R1, R2) sensitivity analysis across multiple $w$; w=10w{=}10 reported as a cross-study reference Calculate WMCC across w (R2) report the WMCC profile, not a single value REPORTING Report all confusion matrix elements (R4) Report CIs (R5) if subsampling / cross-SR Report null outputs (R6) Release open science artifacts (R7) Report full results; cross-study comparability via chance-anchored metrics
Figure 5: Decision tree for LLM-based SR screening evaluation: benchmarking studies (usually retrospective). Dashed elements indicate conditional steps (R5 CIs apply when the benchmark uses subsampling or assesses cross-SR variation, and may be omitted for a single-dataset benchmark that reports the full confusion matrix; R9 applies only when the study design provides training data). Numbers in parentheses refer to the recommendations in the text; R4 appears in both the Compute and Report steps because the same confusion matrix is computed once and then reported in full. Companion: Figure˜6 for deployment.
DEPLOYMENT FOR A SPECIFIC SYSTEMATIC REVIEW (SR) Should this LLM be used to screen my SR? Operational go/no-go decision. STUDY DESIGNMETRIC SELECTION Prospective design / temporal safeguards (R8) Non-LLM baselines (R9) only if the study aims to inform SR practice and labels are available (e.g., SR updates, early validation samples) Predefine Lost Evidence threshold (R3) Compute confusion matrix (R4) Calculate Lost Evidence/Recall (R1) Calculate MCC (R1) COST ASYMMETRY ASSESSMENT Define FN:FP cost ratio w (R1, R2) Calculate WMCC (R2) REPORTING Report all confusion matrix elements (R4) Report CIs (R5) Report null outputs (R6) Release open science artifacts (R7) Lost Evidence (R1) >> threshold (R3)? Escalate to human-only review (R10) Deploy LLM for screening YESNO
Figure 6: Decision tree for LLM-based SR screening evaluation: deployment for a specific SR. Dashed elements indicate conditional steps (R9 applies primarily in SR updates where prior screening provides training data). The decision point checks whether Lost Evidence exceeds the pre-specified threshold (R3), leading to escalation (R10) or deployment. R4 appears in both the Compute and Report steps because the same confusion matrix is computed once and then reported in full. Companion: Figure˜5 for benchmarking.
As a result, we have the following recommendations organized by target audience and themes:
@@
Target Audience: Researchers and Practitioners Metrics and cost-sensitive evaluation   (R1) Standardize reporting on Lost Evidence (Recall), MCC, and Weighted MCC (WMCC) with explicit justification of FN:FP cost ratios, and report only relevant metrics while avoiding Accuracy/PABAK as primary metrics [Both study types] (origin: (P9) in Section˜2.5 & Section˜2.3). (R2) Base comparative conclusions on cost-sensitive analyses that reflect asymmetric misclassification costs, using WMCC to combine chance-correction with cost asymmetry and avoiding over-optimizing Recall alone [Both study types] (origin: (P9) in Section˜2.5). For benchmarking studies where domain-specific cost ratios are not available, report a sensitivity analysis across multiple $w$ values as the primary result; w=10w{=}10 can additionally be reported as a cross-study reference point for comparability with deployment studies. (R3111An asterisk ’*’ means that the recommendation is optional.) When deploying an LLM for a specific SR, predefine acceptable Lost Evidence (minimum Recall) thresholds as guardrails for the review’s risk tolerance and objectives, aligned to review type (e.g., SR, Mapping/Scoping Study, Rapid Reviews) and domain (e.g., healthcare, software engineering) [Deployment only] (origin: Section˜4.1). Reporting and transparency   (R4) Publish complete confusion matrices for every model, dataset, and prompt to enable recomputation of necessary metrics like Lost Evidence, MCC, WMCC, and alternative (e.g., cost-benefit) analyses and future meta-analyses [^7] DC+ revealed a consistent problem with lost evidence across three reviews and 18 LLMs. This important result is only clear because they reported all their confusion matrices. [Both study types] (origin: (P1) in Section˜2.5). (R5) For validation tests based on samples, report uncertainty for each performance metric via confidence intervals and document the estimation method used. When testing LLMs on a random sample of abstracts to decide whether to deploy them on the remaining abstracts, use resampling without replacement to estimate confidence intervals for likely performance on the full dataset [Conditional — see text] (origin: (P3) in Section˜3.2; resampling method proposed by the authors in Section˜3.5 and illustrated in Figure˜4). For benchmarking studies, the deployment-specific resampling guidance (second clause) does not apply, but CIs remain relevant when the benchmark uses subsampling or assesses cross-SR variation (see Section˜3.5). (R6) Quantify LLM output consistency and null or invalid outputs, specify the evaluations rule that treats unclassifiable or referred-back items for fair metric computation [Both study types] (origin: (P4) and (P5) in Section˜3.2). (R7) Release open science artifacts, including prompts, seeds, code, and curated data, to support open science and independent verification [Both study types] (origin: (P2) in Section˜2.5). Study design and validity   (R8) Use designs that prevent training–test overlap: prospective data collection for deployment studies, or documented temporal/contamination safeguards for retrospective benchmarks; explicitly state the leakage risks and the mitigations applied [Both study types] (origin: (P7) & (P8) in Section˜3.2). (R9) Consider non-LLM baselines when the study aims to inform SR practice (e.g., deployment-situated studies, SR updates) and the design provides labels for training baselines. Pure cross-LLM benchmarks and prompting-strategy evaluations satisfy neither condition and need not include non-LLM baselines [Conditional and optional — see text] (origin: (P2) in Section˜2.5). Decision thresholds and operations   (R10) When observed Lost Evidence exceeds the pre-specified threshold, escalate to human review or adjust prompts/models to maintain SR validity [Deployment only] (origin: Section˜4.1).
Summary: Applicability by Study Type The table below summarises which of R1–R10 apply to each study type. Benchmarking studies do not need R3 and R10 because they do not make deployment decisions; R5’s deployment-specific resampling guidance does not apply, but reporting CIs remains relevant when benchmarks use subsampling or assess cross-SR variation. R9 is conditional and optional in both tracks — optional in the strict sense that researchers are advised to consider non-LLM baselines rather than being required to include them. Benchmarking studies (usually retrospective, no deployment decision)   Required: R1 (metrics: Lost Evidence, MCC, WMCC), R2 (cost-sensitive; sensitivity analysis across multiple $w$ values is the primary result, with w=10w{=}10 reported additionally as a cross-study reference point), R4 (full confusion matrices), R6 (consistency & null outputs), R7 (open artefacts), R8 (leakage safeguards). Conditional: R5 (CIs) — the deployment-specific resampling guidance does not apply, but CIs are relevant when benchmarks assess cross-SR variation or use subsampling (see Section˜3.5); R9 (non-LLM baselines) — not required when the benchmark’s sole aim is cross-LLM comparison or prompting-strategy evaluation; consider including baselines only if the benchmark additionally aims to inform operational SR use (e.g., SR updates, reused labelled datasets from which baselines can be trained). Not applicable: R3 (predefined Lost Evidence threshold), R10 (escalation to human review). Deployment for a specific SR (prospective, validation)   Required: R1, R2, R3, R4, R5, R6, R7, R8, R10. Conditional: R9 — consider non-LLM baselines when the study aims to inform SR practice and a labelled sample is available (from prior screening in an SR update, or from early validation in a new SR, cf. [41]).